OpenStudy (anonymous):

The slope of the tangent line to a curve at any point (x, y) on the curve is x divided by y. What is the equation of the curve if (4, 1) is a point on the curve?

OpenStudy (jhannybean):

So does this mean \[\frac{dy}{dx} = \frac{x}{y}\]

OpenStudy (jhannybean):

If so, \[ydy = xdx\]Im kind of new to this.... just attempting.

OpenStudy (jhannybean):

\[\int ydy=\int xdx\]\[\frac{1}{2}y^2=x^2+C\]\[y=\sqrt{2(x^2+C)}~? \]

OpenStudy (anonymous):

the answer only give them as y^2 and x^2

OpenStudy (jhannybean):

I havent taken ordinary differentials yet, so i'm just attempting, I may be wrong. It looked like a fun problem to try.

OpenStudy (amistre64):

1/2 x^2 + C

OpenStudy (anonymous):

no wait sorry i was wrong

OpenStudy (jhannybean):

Ah ok.

OpenStudy (irishboy123):

there's a mistake in the integration -- look at it again......

OpenStudy (jhannybean):

\[y^2=\frac{1}{2}x^2+C\]

OpenStudy (amistre64):

\[\frac12y^2=\frac12x^2+C\] \[y^2-x^2=C\]

OpenStudy (jhannybean):

Oh is that how it works? I see.

OpenStudy (anonymous):

they have x^2-y^2=15 x^2+y^2=15 x+y=15 xy=15

OpenStudy (amistre64):

who is 'they'?

OpenStudy (jhannybean):

So then we can easily find C right? We're given the point (4,1) and we have our equation \[y^2-x^2=C\]\[(1)^2-(4)^2=C\]\[1-16=C\]\[C=-15\]

OpenStudy (amistre64):

(4,1) is read as (x=4, y=1) but in this case its not really that important

OpenStudy (amistre64):

dbl chk y^2 - x^2 = -15 2y y' - 2x = 0 y y' - x = 0 y' =x/y